Please help solve for y
(y-5)/(x+2) = -7/4

Answer by stanbon(75887)

(y-5)/(x+2) = -7/4
Cross-multiply to get:
4y-20 = -7x-14
4y = -7x+6
y = (-7/4)x+(3/2)
--------------
Comment: You cannot "solve" this as you
only have one equation and you have two
variables.
=================
Cheers,
Stan H.